# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ A.18.4. Proof of Proposition 32
Since $\kappa^{\star}\in[\underline{K},\bar{K}]$,
$\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})=$ $\Bigg{[}\sum_{n=1}^{N}\Bigg{(}-\delta_{n}Y_{n}+(1-Y_{n})\delta_{n}\frac{e^{-\kappa^{\star}\delta_{n}}}{1-e^{-\kappa^{\star}\delta_{n}}}\Bigg{)}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}\Big{(}-1+\frac{e^{-\kappa^{\star}\delta_{0}}}{1-e^{-\kappa^{\star}\delta_{0}}}\Big{)}\Bigg{]}$
$=$ $\sum_{n=1}^{N}\frac{-\delta_{n}\bm{1}_{\delta_{n}<+\infty}}{1-e^{-\kappa^{\star}\delta_{n}}}\Big{(}Y_{n}-e^{-\kappa^{\star}\delta_{n}}\Big{)}+\delta_{0}\Big{(}-1+\frac{e^{-\kappa^{\star}\delta_{0}}}{1-e^{-\kappa^{\star}\delta_{0}}}\Big{)}.$
Let $f(\delta;\kappa^{\star})=\frac{-\delta\bm{1}_{\delta_{n}<+\infty}}{1-e^{-\kappa^{\star}\delta}}$, we know that $f(\delta)$ is bounded given $\delta\in[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$ and $\sup_{\delta\in[\underline{\delta},\bar{\delta}]}\left|f(\delta;\kappa^{\star})\right|=\left|f(\bar{\delta};\kappa^{\star})\right|$. Therefore, by Proposition 28, with $\mathbb{P}^{\star}$-probability at least $1-\epsilon$,
$\left|\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})\right|$ $\leq\left|f(\bar{\delta};\kappa^{\star})\right|\sqrt{4N\ln(\tfrac{2}{\epsilon})}+\delta_{0}\Big{|}-1+\frac{e^{-\kappa^{\star}\delta_{0}}}{1-e^{-\kappa^{\star}\delta_{0}}}\Big{|},$
@@
$C=2\left|f(\bar{\delta};\kappa^{\star})\right|,\quad c=\delta_{0}\Big{|}-1+\frac{e^{-\kappa^{\star}\delta_{0}}}{1-e^{-\kappa^{\star}\delta_{0}}}\Big{|}.$
@@ A.18.5. Proof of Theorem 33
By the mean value theorem, there exists $\lambda\in(0,1)$ such that for $\tilde{\kappa}=\lambda\kappa_{N}+(1-\lambda)\kappa^{\star}$,
$0=\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})+\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\tilde{\kappa})(\kappa_{N}-\kappa^{\star}).$
@@
$\left|\kappa_{N}-\kappa^{\star}\right|=\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\tilde{\kappa})\Big{)}^{-1}\big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})\big{|}\leq\Big{(}\inf_{\kappa\geq 0}\big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\big{)}\Big{)}^{-1}\big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})\big{|}.$
By Proposition 30 and Proposition 32, there exists constants $C,c,C^{\prime},c^{\prime}\geq 0$ such that it holds with $\mathbb{P}^{\star}$-probability at least $1-2\epsilon$ for any $\epsilon\geq 0$,
$\left|\kappa_{N}-\kappa^{\star}\right|\leq\frac{C^{\prime}\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}+c^{\prime}}{cN-C\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}}.$
Let $N_{0}=\frac{4C^{2}}{c^{2}}$. Then, if $N\geq N_{0}\ln(\tfrac{2}{\epsilon})$,
$cN-C\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}$ $\geq\frac{cN}{2}+\left(\frac{c}{2}\sqrt{N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}}\right)\sqrt{N}-C\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}=\frac{cN}{2}.$
Therefore, it it holds with $\mathbb{P}^{\star}$-probability at least $1-2\epsilon$ for any $\epsilon\geq 0$,
$\left|\kappa_{N}-\kappa^{\star}\right|\leq\frac{2C^{\prime}}{c}N^{-1/2}\sqrt{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}+\frac{2c^{\prime}}{c}N^{-1}.$
@@
By choosing ${N_{0}}$ to be sufficiently large, we can guarantee that if $N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}$, then $N\geq\tilde{N_{0}}\ln\big{(}\tfrac{2\pi^{2}N^{2}}{3\epsilon}\big{)}$ where $\tilde{N_{0}}$ is a constant given in Theorem 33.
In particular, for such $N$, Theorem 33 holds with $\epsilon_{N}=\tfrac{3\epsilon}{\pi^{2}N^{2}}$. Let $A_{N}$ denote the corresponding event. We can see that
$\mathbb{P}^{\star}\left(\bigcup_{N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}}A_{N}^{c}\right)\leq\sum_{N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}}\mathbb{P}^{\star}(A_{N}^{c})\leq\sum_{N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}}2\left(\frac{3\epsilon}{\pi^{2}N^{2}}\right)\leq\epsilon.$
@@
A.20. Proof of Proposition 34
On the event that Corollary 33.1 holds, we can see that
$\begin{split}X_{N_{T}}&\leq\sum_{n=0}^{\lfloor N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rfloor}(\tau_{n+1}-\tau_{n})|\underline{K}-\bar{K}|^{2}\\ &\qquad+\sum_{n=\lceil N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rceil}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(Cn^{-1/2}\sqrt{\ln\big{(}\tfrac{2n}{\epsilon}\big{)}}+cn^{-1}\right)^{2}\\ &\leq C^{\prime}\sum_{n=0}^{\lfloor N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rfloor}(\tau_{n+1}-\tau_{n})+C\sum_{n=\lceil N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rceil}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\,.\end{split}$
By choosing $C$ to be sufficiently large, i.e. $C\geq C^{\prime}N_{0}$, we have
(59) $\begin{split}X_{N_{T}}&\leq C^{\prime}(\tau_{1}-\tau_{0})+C^{\prime}\sum_{n=1}^{\lfloor N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rfloor}(\tau_{n+1}-\tau_{n})\\ &\quad-C\sum_{n=1}^{\lfloor N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rfloor}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &\leq C^{\prime}(\tau_{1}-\tau_{0})+\sum_{n=1}^{\lfloor N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\rfloor}(\tau_{n+1}-\tau_{n})\left(C^{\prime}-C\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\right)\\ &\quad+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &\leq C(\tau_{1}-\tau_{0})+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\,,\end{split}$
where the last inequality comes from $C^{\prime}-C\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\leq C^{\prime}-C\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}\leq 0$ for any $1\leq n\leq N_{0}\ln(\tfrac{2}{\epsilon})$.
Next we need the following lemma which is proved in [39, Lemma 3.1].
@@ A.20. Proof of Proposition 34
Let $Y_{n}$ be an IID sub-exponential random variable with mean 00 and $(\rho_{n})\subseteq{\mathbb{R}}$. Then there exists a constant $C$ such that for any $\epsilon>0$ and $N\in{\mathbb{N}}$
$\mathbb{P}\left(\left|\sum_{n=1}^{N}Y_{n}\rho_{n}\right|\geq C\ln\big{(}\tfrac{2}{\epsilon}\big{)}\sqrt{\sum_{n=1}^{N}\rho_{n}^{2}}\right)\leq\epsilon.$
By using the above result applying to $\epsilon_{N}\propto\epsilon/N^{2}$ and taking the countable union of the above events, we have
$\mathbb{P}\left(\left|\sum_{n=1}^{N}Y_{n}\rho_{n}\right|\leq C\ln\big{(}\tfrac{2N}{\epsilon}\big{)}\sqrt{\sum_{n=1}^{N}\rho_{n}^{2}}\quad\text{for all}\quad N\in{\mathbb{N}}\right)\geq 1-\epsilon\,.$
Now, we note that $(\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}})$ is sub-exponential with mean 00. Hence, on the event that Corollary 33.1 and Lemma 38 hold with $Y_{n}=\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}}$, we know that with probability at least $1-2\epsilon$, it holds that
(60) $\begin{split}(\tau_{1}-\tau_{0})+&\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &=\frac{1}{\lambda^{+}+\lambda^{-}}\left(1+\sum_{n=1}^{N_{T}}\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\right)\\ &\qquad+\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}})\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)+(\tau_{1}-\tau_{0}-\frac{1}{\lambda^{+}+\lambda^{-}})\\ &\leq\frac{1}{\lambda^{+}+\lambda^{-}}+\frac{1}{\lambda^{+}+\lambda^{-}}\ln N_{T}\ln\big{(}\tfrac{2}{\epsilon}\big{)}+\frac{1}{2(\lambda^{+}+\lambda^{-})}\ln^{2}N_{T}\\ &\qquad+C\ln\big{(}\tfrac{2(N_{T}+1)}{\epsilon}\big{)}\sqrt{1+\sum_{n=1}^{N_{T}}\left(\frac{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}{n}+\frac{\ln n}{n}\right)^{2}}\\ &\leq\frac{1}{\lambda^{+}+\lambda^{-}}+\frac{1}{\lambda^{+}+\lambda^{-}}\ln N_{T}\ln\big{(}\tfrac{2}{\epsilon}\big{)}+\frac{1}{2(\lambda^{+}+\lambda^{-})}\ln^{2}N_{T}\\ &\qquad+C\big{(}\ln(2N_{T})+\ln\big{(}\tfrac{2}{\epsilon}\big{)}\big{)}\sqrt{3+\left(\ln\left(\tfrac{2}{\epsilon}\right)\right)^{2}+2\ln\left(\tfrac{2}{\epsilon}\right)}\\ &\leq\left(\frac{1}{2(\lambda^{+}+\lambda^{-})}+C\sqrt{3}\right)\ln^{2}N_{T}+\left(\frac{1}{\lambda^{+}+\lambda^{-}}+C\right)\ln N_{T}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\\ &\qquad+C(\sqrt{3}+1+\ln 2)\ln^{2}\big{(}\tfrac{2}{\epsilon}\big{)}+\left(\frac{1}{\lambda^{+}+\lambda^{-}}+C\sqrt{3}\ln 2\right)\,,\end{split}$
where the last inequality uses the fact that $\ln N_{T}\leq\ln^{2}N_{T}$ for large $N_{T}$ and $\ln\big{(}\tfrac{2}{\epsilon}\big{)}\leq\ln^{2}\big{(}\tfrac{2}{\epsilon}\big{)}$ for small $\epsilon$. Note that the constant $C$ comes from Lemma 38 which is independent of $\epsilon$, hence the result.

### v2

@@ A.18.4. Proof of Proposition 34
Since $\kappa^{\ast}\in[\underline{K},\bar{K}]$, by (28),
$\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})=$ $\Bigg{[}\sum_{n=1}^{N}\Bigg{(}-\delta_{n}Y_{n}+(1-Y_{n})\delta_{n}\frac{e^{-\kappa^{\ast}\delta_{n}}}{1-e^{-\kappa^{\ast}\delta_{n}}}\Bigg{)}+\delta_{0}\Big{(}-1+\frac{e^{-\kappa^{\ast}\delta_{0}}}{1-e^{-\kappa^{\ast}\delta_{0}}}\Big{)}\Bigg{]}$
$=$ $\sum_{n=1}^{N}\frac{-\delta_{n}\mathds{1}_{\delta_{n}<+\infty}}{1-e^{-\kappa^{\ast}\delta_{n}}}\Big{(}Y_{n}-e^{-\kappa^{\ast}\delta_{n}}\Big{)}+\delta_{0}\Big{(}-1+\frac{e^{-\kappa^{\ast}\delta_{0}}}{1-e^{-\kappa^{\ast}\delta_{0}}}\Big{)}\,,$
where the last equality comes from the fact that $Y_{n}=0$ a.s. when $\delta_{n}=+\infty$ as discussed in Remark 15. Let $f(\delta;\kappa^{\ast})=\frac{-\delta\mathds{1}_{\delta_{n}<+\infty}}{1-e^{-\kappa^{\ast}\delta}}$, we know that $f(\delta)$ is bounded given $\delta\in[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$ and $\sup_{\delta\in[\underline{\delta},\bar{\delta}]}\left|f(\delta;\kappa^{\ast})\right|=\left|f(\bar{\delta};\kappa^{\ast})\right|$. Therefore, by Proposition 31, with $\mathbb{P}^{\ast}$-probability at least $1-\varepsilon$,
$\left|\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})\right|\leq\left|f(\bar{\delta};\kappa^{\ast})\right|\sqrt{4N\ln(\tfrac{2}{\varepsilon})}+\delta_{0}\Big{|}-1+\frac{e^{-\kappa^{\ast}\delta_{0}}}{1-e^{-\kappa^{\ast}\delta_{0}}}\Big{|}\,,$
@@
$C=2\left|f(\bar{\delta};\kappa^{\ast})\right|,\quad c=\delta_{0}\Big{|}-1+\frac{e^{-\kappa^{\ast}\delta_{0}}}{1-e^{-\kappa^{\ast}\delta_{0}}}\Big{|}.$
@@ A.18.5. Proof of Theorem 35
By the mean value theorem, there exists $\lambda\in(0,1)$ such that for $\tilde{\kappa}=\lambda\kappa_{N}+(1-\lambda)\kappa^{\ast}$,
$0=\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})+\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\tilde{\kappa})(\kappa_{N}-\kappa^{\ast})\,.$
@@
$\left|\kappa_{N}-\kappa^{\ast}\right|=\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\tilde{\kappa})\Big{)}^{-1}\big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})\big{|}\leq\Big{(}\inf_{\kappa>0}\big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\big{)}\Big{)}^{-1}\big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})\big{|}\,.$
By Proposition 32 and Proposition 34, there exists constants $C,c,C^{\prime},c^{\prime}\geq 0$ such that it holds with $\mathbb{P}^{\ast}$-probability at least $1-2\varepsilon$ for any $\varepsilon\geq 0$,
$\left|\kappa_{N}-\kappa^{\ast}\right|\leq\frac{C^{\prime}\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}+c^{\prime}}{cN-C\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}}\,.$
Let $N_{0}=\frac{4C^{2}}{c^{2}}$. Then, if $N\geq N_{0}\ln(\tfrac{2}{\varepsilon})$,
$cN-C\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}$ $\geq\frac{cN}{2}+\left(\frac{c}{2}\sqrt{N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}\right)\sqrt{N}-C\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}=\frac{cN}{2}\,.$
Therefore, it holds with $\mathbb{P}^{\ast}$-probability at least $1-2\varepsilon$ such that for any $\varepsilon\geq 0$,
$\left|\kappa_{N}-\kappa^{\ast}\right|\leq\frac{2C^{\prime}}{c}N^{-1/2}\sqrt{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}+\frac{2c^{\prime}}{c}N^{-1}\,.$
@@
By choosing ${N_{0}}$ to be sufficiently large, we can guarantee that if $N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}$, then $N\geq\tilde{N_{0}}\ln\big{(}\tfrac{2\pi^{2}N^{2}}{3\varepsilon}\big{)}$ where $\tilde{N_{0}}$ is a constant given in Theorem 35.
In particular, for such $N$, Theorem 35 holds with $\varepsilon_{N}=\tfrac{3\varepsilon}{\pi^{2}N^{2}}$. Let $A_{N}$ denote the corresponding event. We can see that
$\mathbb{P}^{\ast}\Bigg{(}\bigcup_{N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}A_{N}^{c}\Bigg{)}\leq\sum_{N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}\mathbb{P}^{\ast}(A_{N}^{c})\leq\sum_{N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}2\left(\frac{3\varepsilon}{\pi^{2}N^{2}}\right)\leq\varepsilon.$
@@
A.20. Proof of Corollary 35.2
Proof.
Let $C,c,N_{0}\geq 0$ be the constants from Corollary 35.1 and $N_{0}^{\prime}=(\tfrac{c}{C})^{2}$, then it holds with $\mathbb{P}^{\ast}$-probability at least $1-\varepsilon$ such that for any $\varepsilon\geq 0$,
$\begin{split}\left|\kappa_{N}-\kappa^{\ast}\right|&\leq CN^{-1/2}\sqrt{\ln(\tfrac{2N}{\varepsilon})}+cN^{-1}\\ &\leq 2CN^{-1/2}\sqrt{\ln(\tfrac{2N}{\varepsilon})}\qquad\text{for all }N\geq\max\left(N_{0}\ln(\tfrac{2}{\varepsilon}),N_{0}^{\prime}/\ln(\tfrac{2}{\varepsilon})\right)\,.\end{split}$
For such $N$, we have
$\kappa_{N}\leq\kappa^{\ast}+2CN^{-1/2}\sqrt{\ln(\tfrac{2N}{\varepsilon})}\,,$
and
$\kappa_{N}\geq\kappa^{\ast}-2CN^{-1/2}\sqrt{\ln(\tfrac{2N}{\varepsilon})}\,.$
Let $N_{1}^{\prime}=\max\left(\bar{K}-\kappa^{\ast},\kappa^{\ast}-\underline{K}\right)$ and $N_{1}=(\tfrac{N_{1}^{\prime}}{2C})^{2}$. Then it holds with $\mathbb{P}^{\ast}$-probability at least $1-\varepsilon$ such that for any $\varepsilon\geq 0$,
$\kappa_{N}\in[\underline{K},\bar{K}]\quad\text{ for all }N\geq\max\left(\ln(\tfrac{2}{\varepsilon})/(N_{1}-1),N_{0}\ln(\tfrac{2}{\varepsilon}),N_{0}^{\prime}/\ln(\tfrac{2}{\varepsilon})\right)\,,$
which completes the proof. ∎
A.21. Proof of Proposition 36
On the event that Corollary 35.1 and Corollary 35.2 hold, we can see that
$\begin{split}X_{N_{T}}&\leq\sum_{n=0}^{\lfloor N^{\prime}\rfloor}(\tau_{n+1}-\tau_{n})|\underline{K}-\bar{K}|^{2}\\ &\qquad+\sum_{n=\lceil N^{\prime}\rceil}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(Cn^{-1/2}\sqrt{\ln\big{(}\tfrac{2n}{\varepsilon}\big{)}}+cn^{-1}\right)^{2}\\ &\leq C^{\prime}\sum_{n=0}^{\lfloor N^{\prime}\rfloor}(\tau_{n+1}-\tau_{n})+C\sum_{n=\lceil N^{\prime}\rceil}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\,,\end{split}$
where $N^{\prime}=\max\Big{(}N_{0}\ln(\tfrac{2}{\varepsilon}),N_{0}^{\prime}/\ln(\tfrac{2}{\varepsilon})\Big{)}$. By choosing $C$ to be sufficiently large, we have
$\begin{split}X_{N_{T}}&\leq C^{\prime}(\tau_{1}-\tau_{0})+C^{\prime}\sum_{n=1}^{\lfloor N^{\prime}\rfloor}(\tau_{n+1}-\tau_{n})\\ &\quad-C\sum_{n=1}^{\lfloor N^{\prime}\rfloor}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &\leq C^{\prime}(\tau_{1}-\tau_{0})+\sum_{n=1}^{\lfloor N^{\prime}\rfloor}(\tau_{n+1}-\tau_{n})\left(C^{\prime}-C\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\right)\\ &\quad+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &\leq C(\tau_{1}-\tau_{0})+C\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\,,\end{split}$
where the last inequality comes from $C^{\prime}-C\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\leq C^{\prime}-C\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}\leq 0$ for any $1\leq n\leq N^{\prime}$ under a sufficiently large $C$.
Next we need the following lemma which is proved in [40, Lemma 3.1].
@@ A.21. Proof of Proposition 36
Let $Y_{n}$ be an IID sub-exponential random variable with mean 00 and $(\rho_{n})\subseteq{\mathbb{R}}$. Then there exists a constant $C$ such that for any $\varepsilon>0$ and $N\in{\mathbb{N}}$
$\mathbb{P}\left(\left|\sum_{n=1}^{N}Y_{n}\rho_{n}\right|\geq C\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\sqrt{\sum_{n=1}^{N}\rho_{n}^{2}}\right)\leq\varepsilon.$
By using the above result applying to $\varepsilon_{N}\propto\varepsilon/N^{2}$ and taking the countable union of the above events, we have
$\mathbb{P}\left(\left|\sum_{n=1}^{N}Y_{n}\rho_{n}\right|\leq C\ln\big{(}\tfrac{2N}{\varepsilon}\big{)}\sqrt{\sum_{n=1}^{N}\rho_{n}^{2}}\quad\text{for all}\quad N\in{\mathbb{N}}\right)\geq 1-\varepsilon\,.$
Now, we note that $(\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}})$ is sub-exponential with mean 00. Hence, on the event that Corollary 35.1 and Lemma 40 hold with $Y_{n}=\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}}$, we know that with probability at least $1-2\varepsilon$, it holds that
(63) $\begin{split}(\tau_{1}-\tau_{0})+&\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\\ &=\frac{1}{\lambda^{+}+\lambda^{-}}\left(1+\sum_{n=1}^{N_{T}}\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)\right)\\ &\qquad+\sum_{n=1}^{N_{T}}(\tau_{n+1}-\tau_{n}-\frac{1}{\lambda^{+}+\lambda^{-}})\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)+(\tau_{1}-\tau_{0}-\frac{1}{\lambda^{+}+\lambda^{-}})\\ &\leq\frac{1}{\lambda^{+}+\lambda^{-}}+\frac{1}{\lambda^{+}+\lambda^{-}}\ln N_{T}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}+\frac{1}{2(\lambda^{+}+\lambda^{-})}\ln^{2}N_{T}\\ &\qquad+C\ln\big{(}\tfrac{2(N_{T}+1)}{\varepsilon}\big{)}\sqrt{1+\sum_{n=1}^{N_{T}}\left(\frac{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}{n}+\frac{\ln n}{n}\right)^{2}}\\ &\leq\frac{1}{\lambda^{+}+\lambda^{-}}+\frac{1}{\lambda^{+}+\lambda^{-}}\ln N_{T}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}+\frac{1}{2(\lambda^{+}+\lambda^{-})}\ln^{2}N_{T}\\ &\qquad+C\big{(}\ln(2N_{T})+\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\big{)}\sqrt{3+\left(\ln\left(\tfrac{2}{\varepsilon}\right)\right)^{2}+2\ln\left(\tfrac{2}{\varepsilon}\right)}\\ &\leq\left(\frac{1}{2(\lambda^{+}+\lambda^{-})}+C\sqrt{3}\right)\ln^{2}N_{T}+\left(\frac{1}{\lambda^{+}+\lambda^{-}}+C\right)\ln N_{T}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\\ &\qquad+C(\sqrt{3}+1+\ln 2)\ln^{2}\big{(}\tfrac{2}{\varepsilon}\big{)}+\left(\frac{1}{\lambda^{+}+\lambda^{-}}+C\sqrt{3}\ln 2\right)\,,\end{split}$
where the last inequality uses the fact that $\ln N_{T}\leq\ln^{2}N_{T}$ for large $N_{T}$ and $\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\leq\ln^{2}\big{(}\tfrac{2}{\varepsilon}\big{)}$ for small $\varepsilon$. Note that the constant $C$ comes from Lemma 40 which is independent of $\varepsilon$, hence the result.